# Augmenting LLMs to Repair Obsolete Test Cases with Static Collector and Neural Reranker

### v1

@@ Abstract
During software evolution, it is advocated that test code should co-evolve with production code. In real development scenarios, test updating may lag behind production code changing, which may cause the project to fail to compile or bring other troubles. Existing techniques based on pre-trained language models can be adopted to repair obsolete tests caused by such unsynchronized code changes, especially syntactic-related ones. However, the lack of target-oriented contextual information affects repair accuracy on large-scale projects. Starting from an obsoleted test, the key challenging task is precisely identifying and constructing Test-Repair-Oriented Contexts (TROCtx) from the whole repository within a limited token size.
In this paper, we propose Synbciatr (Syntactic-Breaking-Change-Induced Automated Test Repair), a novel approach to automatically repair obsolete test cases via precise and concise TROCtx construction. Inspired by developers’ programming practices of the task, we design three types of TROCtx: class contexts, usage contexts, and environment contexts. For every type of TROCtx, Synbciatr automatically collects the changed-token-related code information through static analysis techniques. Then it generates reranking queries to identify the most relevant TROCtxs, which will be taken as the repair-required key context and be input to Large Language Model for the final test repair.
To evaluate the effectiveness of Synbciatr, we construct a benchmark dataset that contains diverse syntactic breaking changes. The experimental results show that Synbciatr outperforms baseline approaches both on textual- and intent-matching metrics. With the augmentation of TROCtx constructed by Synbciatr, hallucinations are reduced by 57.1%.
@@
Software Evolution, Test Repair, LLM, Static AnalysisI Introduction
Software evolution is a fundamental and significant aspect of software development [1]. For large-scale software such as Kafka [2], the project frequently evolves, where a new version is released in one to two weeks on average and new commits are submitted nearly every day. As production code usually changes during software evolution, it is crucial to maintain and co-evolve associated test codes to ensure that they remain effective in validating the software’s functionality [3], [4]. Specifically, for software that is invoked as libraries, it is important to co-evolve the test code to follow the changes of production code, which can help developers quickly notice the backward incompatible changes that may affect its clients [5].
To automatically co-evolve production code and test code, previous studies analyze and mine the software code to extract production-test co-evolution rules and patterns [6], [7]. However, as real-world code changes come in a great many forms, they are hard to summarize into a small number of general patterns. In recent years, with the rapid advancement of machine learning and large language models (LLM) [8], [9], many studies have utilized learning-based techniques to assist the production and test code co-evolution, including obsolete test cases identification [10], production-test co-evolution pairs extraction [11], etc., which yielded favorable results. To further reduce the developer’s burden, researchers also try to repair the obsolete test cases automatically. For example, Hu et al. [12] identify and update obsolete test methods by fine-tuning a pre-trained model initialized from CodeT5 [13], which is currently the SOTA approach for repairing obsolete test cases.
Though the learning-based techniques resulted in positive outcomes on many tasks, they still faced with difficulties when operating code in complex repositories. For example, API signature changing is the most common and straightforward code-changing type in software evolution. However, focusing on this type of production change, the SOTA work Ceprot [12] failed to correctly repair about three-quarters of test cases on real-world projects (refer to Section V-A), which means there still have a gap before it can be applied to co-evolve tests in practice.
To fill in the gap, we mainly target precisely repairing these signature-related code changes based on the power of LLMs. Here, the signature-based code changes are also called Syntactic Breaking Changes (SynBC) as they may lead to compilation errors if their associated tests are not co-evolved. According to our investigation, accurately repairing synBC-related obsolete test cases based on LLMs still faces the following key challenges.
@@
The repair-oriented code contexts are unclear. When developers manually repair a test case, not only the changed signature of the focal method but also many other changed-token-related code contexts are required. When fixing tests with learning-based approaches, the repair-oriented code context must be explicitly extracted. As existing works repair test methods solely with the original and updated focal methods as input, it is still unclear which type of contexts are essential for accurate repair and how to extract them.
@@
The token size of code contexts is limited. Compared to small-scale models, LLMs have demonstrated their extraordinary capabilities. However, even though they are designed with increasingly larger context windows, it is impractical to simply include the entire repository as input. And extra irrelevant information may bring negative effects as well. That means the extracted code context is not the more the better. As the number of context tokens processed in a single interaction is restricted, for all the code contexts that have a relation with the changed signature, it is required to sort and pick out the most relevant ones as part of the final context.
To this end, we propose Synbciatr (Syntactic-Breaking-Changes-Induced Automated Test Repair), a novel approach to repairing obsolete test cases caused by SynBC at the method level. To address challenge C1, Synbciatr designs three types of contexts, including Class Contexts (ClassCtx), Usage Contexts (UsgCtx), and Environment Contexts (EnvCtx) as Test-Repair-Oriented Contexts (TROCtx). These contexts focus on different aspects of changed code, but all aim at providing adequate contextual information for test-repairing together. For challenge C2, Synbciatr identifies and constructs qualified TROCtx from the repository in two stages. ❶ First, Synbciatr collects all types of TROCtx by static code analysis, specifically, by using Language Server. ❷ Then, inspired by the idea of retrieval-augmented generation (RAG) [14], Synbciatr generates reranking queries to identify the most relevant TROCtxs according to the repair requirement extracted from the original test case. That process is based on Neural Rerankers [15], a set of neural models for reranking in information retrieval. After constructing TROCtxs, Synbciatr aggregates test-repair-required information to a full LLM prompt and finally generates the repaired test cases.
To evaluate the effectiveness of Synbciatr, we construct a benchmark dataset based on existing work which consists of 136 samples with diverse SynBCs. The evaluation is based on both the textual match and intent match metrics. In terms of the textual match, Synbciatr achieves the best performance against baselines specifically on CodeBLEU (83.3), DiffBLEU (46.7), and Accuracy (32.4). In terms of the intent match, we conduct a human evaluation on verifying whether test cases are correctly repaired without changing their original intent, in which Synbciatr achieve improvements of 248.6% and 9.8% when compared to Ceprot and naiveLLM respectively. Moreover, Synbciatr is capable of reducing 57.1% hallucinations.
@@
We design three types of TROCtx to provide adequate contextual information for test repair caused by SynBC.
We propose Synbciatr to construct TROCtx by combining static collector and neural rerank, and use it to repair test based on LLM.
Experimental results on the benchmark dataset demonstrate that Synbciatr can repair obsolete test cases caused by SynBC more effectively compared to both Ceprot and Naivellm.
The data and code are both publicly available at: https://anonymous.4open.science/r/SynBCIATR-3DD2.
@@ II-A Task Definition
Referring to previous studies [16], [17], [5], Syntactic Breaking issues (SynB) represent signature-based errors in API evolution, such as ClassNotFoundException and NoSuchMethodError. In this paper, we name method changes that may cause SynB as Syntactic Breaking Changes (SynBC), in which the signature of the method changes during evolution.
We define the signature of a method (Method Signature) as a tuple , where:
@@
Given a method that changes from to , we define Syntactic Breaking Changes (SynBC) as:
where and denote the method signatures before and after the change respectively. To figure out whether SynBCs are common in production-test co-evolution, we conducted an empirical study to count the SynBCs exist in the dataset released by Ceprot [12]. Each sample in the dataset contains a change of focal method on a real-world commit from GitHub. It reveals that signature-based focal changes occur in over 40% samples of the dataset, in which parameter and return type-related changes account for a large proportion.
Thus, we further categorize SynBC into the following three types according to the changed elements in a method signature. Note that, a SynBC whose parameter and return types in the method signature both changes belongs to ParamSynBC and RetSynBC at the same time.
@@
Normal Syntactic Changes (NormSynBC). For a SynBC from to , if it does not belong to ParamSynBC or RetSynBC, it is a NormSynBC.
As repairing tests caused by SynBC is a primary work for effective production-test co-evolution, we make efforts to repair obsolete test cases under this type. Here, for a SynBC that happens between the original focal method and the updated focal method , denotes the repository code that method belongs to, and denotes the obsolete test case associated with . Based on this knowledge, we hope to get the updated test case as the final output. The whole task consists of two main steps. First, we extract the repair-oriented contexts for test from repository according to the SynBC from to . That process can be denoted as:
where represents the constructed contexts. Then, with the contexts , repairing obsolete tests can be expressed as:
@@ II-B Language Server and Neural Reranker
Language Server consists of programming language specific tools like static analyzers and compilers. In modern Integrated Development Environments (IDEs), Language Server provides language-specific features like ‘autocomplete’, ‘goto definition’, ‘find usages’, and others [18]. To help IDEs communicate with Language Server [19], Microsoft has created a standard JSON-RPC-based protocol, called Language Server Protocol (LSP), based on which multiple IDEs can communicate with the same Language Server to access intelligence features.
Neural Reranker is a type of machine learning models [20] used to reorder a given set of documents based on their relevance to the query, which is the initial request for information expressed as keywords or complex expressions [15]. It is widely used in the context of information retrieval, natural language processing, and recommendation systems. Instead of ranking based on simple heuristics like the frequency of query terms appearing in the query, Neural Rerankers are trained to take into account more complex features, like the semantic similarity between the query and the document.
@@ II-C Motivating Example
The example in Fig. 1 is used to demonstrate our motivation. It was collected from a real commit in project Alluxio [21]. As shown in the given commit, there is a ParamSynBC for the focal method (named mount), where the third parameter changes from MountOptions to MountPOptions. If the associated original test case (also named mount) does not co-evolve with the change of focal method, the test will fail to be built with compilation error (cannot resolve type).
To automatically repair the obsolete test method, we first apply existing learning-based techniques directly. However, both Ceprot and GPT-4 fail due to using an undefined method (MountPOptions.defaults()) to construct the third parameter, where the correct method invocation should be MountPOptions.getDefaultInstance(). This hallucination occurs for lacking test-repair-oriented contexts (TROCtx) in the input causes. After providing the required contexts shown in Fig. 1, either including the definition of getDefaultInstance in class MountPOptions or the usage change of focal method in other production code, LLMs can generate the correct repaired test.
When developers are asked to repair test for this case, it is convenient for them to refer to the related contexts (TROCtx in Fig. 1) with the programming features (such as ‘goto definition’ and ‘find usages’) provided by Language Servers in IDEs. Inspired by the practices, Synbciatr constructs TROCtx by simulating developers’ behaviors in IDEs. Specifically, Synbciatr collects related contexts by directly interacting with Language Server, and filters out unnecessary ones based on Neural Rerankers. Finally, Synbciatr uses LLMs to repair obsolete test cases with constructed TROCtx.
@@ III Methodology
In this section, we first introduce the overall framework of Synbciatr. Then we demonstrate all the types of TROCtx identified by Synbciatr. Finally, we will describe the main modules in Synbciatr in detail.
@@ III-A Overview
The overall pipeline of Synbciatr is depicted in Fig. 2. Given the change of focal method and the obsolete test method as inputs, Synbciatr consists of three major steps. (1) Collecting TROCtx: Synbciatr analyzes all the related contexts from inputs and requests Language Server to collect and process them into candidate chunks; (2) Reranking TROCtx: Synbciatr reranks candidate chunks with queries constructed from the inputs; (3) Generating full prompt: Synbciatr aggregates the inputs and final TROCtx to generate the full prompt, which is used to repair test with LLM. We also provide a detailed overview of Synbciatr in Fig. 3.
@@ III-B Types of TROCtx
Based on the characteristics of the test-repair task, we categorize TROCtx into three types: class contexts (ClassCtx), usage contexts (UsageCtx), and environment contexts (EnvCtx). In alignment with real-world developers’ practices, these categories of contexts can provide comprehensive and sufficient contextual information for test repair.
@@
ClassCtx: the member accesses of a specific class type, including method accesses and field accesses. These contexts indicate the accurate operations supported by a given class type, serving to alleviate the hallucination of LLMs. Since new class types are introduced in parameter types and the return type, ClassCtx will be collected specially for ParamSynBC and RetSynBC. For example, the ClassCtx for the case caused by ParamSynBC in Fig. 1 is partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // methods are simplified as signatures … public static final int READONLY_FIELD_NUMBER = 1; public boolean hasReadOnly(); public static MountPOptions getDefaultInstance(); … ClassCtx of the new class MountPOptions.
@@
UsageCtx: the changes of usages for the focal method. Usages of the updated focal method in other parts of the repository illustrate how to correctly call the method in the associated test. UsageCtx will be collected for all the types of SynBC. For example, the UsageCtx for case introduced in Fig. 1 is partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff … - mount(alluxioPath, ufsPath, MountOptions.defaults() + mount(alluxioPath, ufsPath, MountPOptions.getDefaultInstance() … UsageCtx of the focal method mount().
@@
EnvCtx: the environmental changes of the focal method and the test method. Given a method , we define the class containing it and its parent classes as the environment of . These contexts indicate modifications of related identifiers and similar change patterns that exist externally to the method. EnvCtx will be collected for all types of SynBC. For example, the EnvCtx of the focal method for the case introduced in Fig. 1 is partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff … - return openFile(path, OpenFileOptions.defaults()); + return openFile(path, OpenFileOptions.getDefaultInstance()); … EnvCtx of of the focal method mount().
@@
Synbciatr aims to collect and rerank different types of TROCtx respectively. In the following subsections, we will describe the details of how each type of TROCtx is collected, reranked, and ultimately aggregated as summarized in Fig. 3.
III-C Collecting TROCtx

### v2

# Fix the Tests: Augmenting LLMs to Repair Test Cases with Static Collector and Neural Reranker

@@ Abstract
During software evolution, it is advocated that test code should co-evolve with production code. In real development scenarios, test updating may lag behind production code changing, which may cause compilation failure or bring other troubles. Existing techniques based on pre-trained language models can be directly adopted to repair obsolete tests caused by such unsynchronized code changes, especially syntactic-related ones. However, the lack of task-oriented contextual information affects the repair accuracy on large-scale projects. Starting from an obsolete test, the key challenging task is precisely identifying and constructing Test-Repair-Oriented Contexts (TROCtxs) from the whole repository within a limited token size.
In this paper, we propose Synter (SYNtactic-breaking-changes-induced TEst Repair), a novel approach based on LLMs to automatically repair obsolete test cases via precise and concise TROCtxs construction. Inspired by developers’ programming practices, we design three types of TROCtx: class context, usage context, and environment context. Given an obsolete test case to repair, Synter firstly collects the related code information for each type of TROCtx through static analysis techniques automatically. Then, it generates reranking queries to identify the most relevant TROCtxs, which will be taken as the repair-required key contexts and be input to the large language model for the final test repair.
To evaluate the effectiveness of Synter, we construct a benchmark dataset that contains a set of obsolete tests caused by syntactic breaking changes. The experimental results show that Synter outperforms baseline approaches both on textual- and intent-matching metrics. With the augmentation of constructed TROCtxs, hallucinations are reduced by 57.1%.
@@
Software Evolution, Obsolete Test Repair, LLM, Static AnalysisI Introduction
Software evolution is a fundamental and significant aspect of software development [1]. For large-scale software such as Kafka [2], the project frequently evolves, where a new version is released in one to two weeks on average and new commits are submitted nearly every day. As production code usually changes during software evolution, it is crucial to maintain and co-evolve associated test code to ensure that they remain effective in validating the software’s functionality [3], [4]. Specifically, for software invoked as libraries, it is important to co-evolve the test code to follow the production code changes, which can help developers quickly notice the backward incompatible changes that may affect its clients [5].
To automatically co-evolve production code and test code, previous studies analyze and mine the software code to extract production-test co-evolution rules and patterns [6], [7]. However, as real-world code changes come in a great many forms, they are hard to summarize into a small number of general patterns. In recent years, with the rapid advancement of machine learning and Large Language Models (LLMs) [8], [9], [10], many studies have utilized learning-based techniques to assist the production and test code co-evolution, including obsolete test case identification [11], production-test co-evolution pair extraction [12], etc., which yielded favorable results. To further reduce the developer’s burden, researchers are also concerned with repairing obsolete test cases automatically. For example, Hu et al. [13] identified and updated obsolete test methods by fine-tuning a pre-trained model initialized from CodeT5 [14], which is currently the SOTA approach for repairing obsolete test cases.
For this task, though directly using learning-based techniques resulted in some positive outcomes, it still faces difficulties in complex repositories. For example, API signature changing is the most common and straightforward code-changing type during software evolution. However, focusing on this type of production change, the SOTA work Ceprot [13] fails to repair about three-quarters of test cases on real-world projects (refer to Table II in Section V-A), which means there is still some gap before existing learning-based approaches can be applied in practice.
To fill this gap, we target repairing signature-related code changes using the power of LLMs. Here, the signature-based code changes are also called Syntactic Breaking Changes (SynBCs) as they may lead to compilation errors if associated tests are not co-evolved. According to our investigation, accurately repairing synBC-related obsolete test cases based on LLMs faces the following key challenges.
@@
The repair-oriented code contexts are unclear. When developers manually repair a test case, not only the signature changes of the focal method but also many other related code contexts are considered for better understanding. When fixing tests with learning-based approaches, the repair-oriented code context must be explicitly extracted. As existing works repair test cases solely with the original and updated focal methods, they are unaware of related contexts. It is essential to determine which types of contexts are essential for accurate repair and how to extract them.
@@
The token size of code contexts is limited. Compared to small-scale models, LLMs have demonstrated their extraordinary capabilities. However, even though they are designed with increasingly larger context windows, it is impractical to simply include the entire repository contents as input. Moreover, extra irrelevant information may bring negative effects as well, which means that the extracted code context is not the more the better. As the number of context tokens for LLM input is restricted, for all the code contexts that may have relations with the changed signature, it is required to sort and pick out the most relevant ones as input.
To this end, we propose Synter (SYNtactic-breaking-changes-induced TEst Repair, originally called Synbciatr), a novel approach for repairing obsolete test cases caused by SynBCs at the method level. To address challenge C1, Synter designs three types of contexts, including Class Contexts (ClassCtxs), Usage Contexts (UsageCtxs), and Environment Contexts (EnvCtxs) as Test-Repair-Oriented Contexts (TROCtxs). These contexts focus on different aspects of changed codes to provide adequate contextual information for test repair. For challenge C2, Synter identifies and constructs qualified TROCtx from the repository in two stages. ❶ First, Synter collects all types of TROCtx by static code analysis, specifically, using the Language Server [15]. ❷ Then, inspired by the idea of Retrieval-Augmented Generation (RAG) [16], Synter generates reranking queries to identify the most relevant TROCtxs in each type according to the repair requirement extracted from the original test case, which is based on Neural Rerankers [17]. After constructing TROCtxs, Synter aggregates the test-repair-required information to a final prompt and generates the repaired test case.
To evaluate the effectiveness of Synter, we construct a benchmark dataset based on existing work which consists of 136 samples with diverse SynBCs. The evaluation is based on both the textual match and intent match metrics. In terms of the textual match, Synter achieves the best performance against baselines specifically on CodeBLEU (83.3), DiffBLEU (46.7), and Accuracy (32.4%). In terms of the intent match, we conduct a human evaluation on verifying whether test cases are correctly repaired without changing their original intents, in which Synter correctly repairs 90.4% test cases, achieving improvements of 248.6% and 9.8% when compared to Ceprot and Naivellm respectively. Moreover, Synter is capable of reducing 57.1% hallucinations caused by Naivellm.
@@
We design three types of TROCtx to provide adequate contextual information for repairing obsolete test cases caused by SynBCs.
We propose Synter to construct TROCtx by combining the static collector and neural reranker, which is utilized to enhance the repairing ability of naive LLM.
Experimental results on the benchmark dataset demonstrate that Synter can repair obsolete test cases caused by SynBCs more effectively compared to both Ceprot and Naivellm.
The data and code are both publicly available at: https://github.com/nonsense-j/SynTeR.
@@ II-A Task Definition
Referring to previous studies [5], [18], [19], Syntactic Breaking (SynB) issues represent signature-based compilation errors in API evolution, such as ClassNotFoundException and NoSuchMethodError. In this paper, we name method signature changes that may cause SynB issues as Syntactic Breaking Changes (SynBCs).
We define the signature of a method (Method Signature) as a 5-tuple , where:
@@
Given a method that changes from to , we define a Syntactic Breaking Change (SynBC) as:
where and denote the method signatures before and after the change respectively. To determine whether SynBCs are common in production-test co-evolution, we conducted an empirical study to count the SynBCs in the breaking change dataset released by Ceprot [13]. Each sample in the dataset contains a change of focal method on a real-world commit from GitHub. It reveals that signature-based focal changes occur in over 40% samples of the dataset, in which parameter and return type-related changes account for a large proportion.
Specifically, we further categorize SynBCs into the following three types according to the changed elements in the method signature. Note that, a SynBC whose parameter and return types in the method signature both change is a ParamSynBC and RetSynBC at the same time.
@@
Normal Syntactic Change (NormSynBC). For a SynBC from to , if it does not belong to ParamSynBC or RetSynBC type, it is a NormSynBC.
Here, for a SynBC that happens between the original focal method and the updated focal method , denotes the repository code that the method belongs to, and denotes the obsolete test case associated with . Based on this knowledge, the test-repair task hopes to get the repaired test case as the final output. The whole task consists of two main steps. First, we extract the repair-oriented contexts for test from repository according to the SynBC from to . That process can be denoted as:
where represents the constructed contexts. Then, we repair the obsolete test with the constructed contexts , which can be expressed as:
@@ II-B Language Server and Neural Reranker
A Language Server consists of programming language-specific tools like static analyzers and compilers. In modern Integrated Development Environments (IDEs), language servers provide language-specific features like ‘autocomplete’, ‘goto definition’, ‘find usages’, and others [20]. Recently, the Microsoft team has created a standard JSON-RPC-based protocol, called Language Server Protocol (LSP) [15], based on which multiple IDEs can communicate with the same language server to access intelligent programming features.
A Neural Reranker is a type of machine learning model [21] used to reorder a given set of documents based on their relevance to a given query, which is the initial request for information expressed as keywords or complex expressions [17]. It is widely used in research fields such as information retrieval, natural language processing, and recommendation systems. Instead of ranking based on simple heuristics like the frequency of query terms appearing in the query, neural rerankers are trained to take into account more complex features, like the semantic similarity between the query and the documents.
@@ II-C Motivating Example
The example in Fig. 1 is used to demonstrate our motivation. It was collected from a real commit in project Alluxio [22]. As shown in the given commit, there is a ParamSynBC for the focal method (named mount), where the third parameter changes from MountOptions to MountPOptions. If the associated original test case (also named mount) does not co-evolve with the change of the focal method, the test will fail for compilation errors (cannot resolve type).
To automatically repair the obsolete test method, we first apply existing learning-based techniques directly. However, both Ceprot and GPT-4 fail due to using an undefined method (MountPOptions.defaults()) to construct the third parameter, where the correct method invocation should be MountPOptions.getDefaultInstance(). This hallucination occurs for lacking Test-Repair-Oriented Contexts (TROCtxs) in the input. After providing the required contexts shown in Fig. 1, by including either the definition of getDefaultInstance in class MountPOptions or the usage change of focal method in other production code, LLMs can generate the correct repaired test.
When developers are asked to repair test for this case in IDEs, it is convenient for them to refer to the related contexts (TROCtxs in Fig. 1) with the programming features (such as ‘goto definition’ and ‘find usages’) provided by language servers. Inspired by these practices, Synter constructs TROCtxs by simulating developers’ behaviors in IDEs. Specifically, Synter collects related contexts by interacting with the language server and filters out unnecessary ones based on neural rerankers. Finally, Synter uses LLMs to repair obsolete test cases with constructed TROCtxs.
@@ III Methodology
In this section, we first show the overall framework of Synter. Then we demonstrate all the types of TROCtxs identified by Synter. Finally, we introduce the technical details of the main modules in Synter.
@@ III-A Overview
The overall pipeline of Synter is depicted in Fig. 2. Given the change of the focal method and the obsolete test method as inputs, Synter consists of three major steps. (1) Collecting TROCtxs: Synter analyzes all the related contexts from inputs and requests the language server to collect and process them into candidate chunks; (2) Reranking TROCtxs: Synter reranks candidate chunks with queries constructed from the inputs; (3) Generating full prompt: Synter aggregates the inputs and final TROCtxs to generate the full prompt, which is used to repair the test with LLM. We also provide a detailed overview in Fig. 3.
@@ III-B Types of TROCtx
Based on the characteristics of the test-repair task, we categorize TROCtxs into three types: Class Contexts (ClassCtxs), Usage Contexts (UsageCtxs), and Environment Contexts (EnvCtxs). In alignment with real-world developers’ practices, these categories of contexts can provide comprehensive and sufficient contextual information for test repair.
@@
ClassCtxs include the member accesses (method and field accesses) of a specific class and its parent classes. These contexts indicate the accurate operations supported by a given class type, serving to alleviate the hallucination of LLMs. Since new class types can be introduced in parameter types and the return type, ClassCtxs will be collected specially for ParamSynBCs and RetSynBCs. For example, the ClassCtxs for the case caused by a ParamSynBC in Fig. 1 are partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // defined in MountPOptions and its parent classes // methods are simplified as signatures … public static final int READONLY_FIELD_NUMBER = 1; public boolean hasReadOnly(); public static MountPOptions getDefaultInstance(); … ClassCtxs of the new class MountPOptions.
@@
UsageCtxs include the changes of usages for the focal method in the diff format. Usages of the updated focal method in other parts of the repository can illustrate how to properly call the method in the associated test. UsageCtxs will be collected for all the types of SynBCs. For example, the UsageCtxs for case introduced in Fig. 1 are partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff … - mount(alluxioPath, ufsPath, MountOptions.defaults() + mount(alluxioPath, ufsPath, MountPOptions.getDefaultInstance() … UsageCtxs of the focal method mount().
@@
EnvCtxs include the environmental changes of the focal and test method in the diff format. Given a method , we define the class containing it and its parent classes as the environment of . Only code changes external to in the environment will be collected to construct its EnvCtxs. These contexts can indicate updates of related identifiers and similar change patterns. EnvCtxs will be collected for all the types of SynBCs. For example, the EnvCtxs of the focal method for the case introduced in Fig. 1 are partly demonstrated in Listing ‣ III-B.
@@
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff // code diffs of class containing mount and its parents … - return openFile(path, OpenFileOptions.defaults()); + return openFile(path, OpenFileOptions.getDefaultInstance()); … EnvCtxs of of the focal method mount().
@@
Synter aims to collect and rerank different types of TROCtxs respectively. In the following subsections, we will describe the details of how each type of TROCtx is collected, reranked, and ultimately aggregated as displayed in Fig. 3.
III-C Collecting TROCtxs